# name=Sara Hooker

### v1

@@ 3.1 Challenges of using FLOP as a metric
FLOP only accounts for a single model, but does not capture risk of the overall system. The emphasis on compute thresholds as an indicator of risk also implies that risk is the property of a single model rather than the system in which it is deployed. In the real-world, impact and risk are rarely attributable to a single model but are a facet of the entire system a model sits in and the way it interacts with its environment (Zaharia et al., 2024; Sculley et al., 2015; Jatho et al., 2023; Raji et al., 2020). Many real-world production systems are made up of cascading models where the final output is produced as a results of inputs being processed by multiple algorithms in sequence (Paleyes et al., 2022; Forum, 2023; Sculley et al., 2015; Shankar et al., 2022). There has yet to be guidance on whether the FLOP threshold is specific to a single model or whether all models that constitute an end-to-end system contribute to the final tally. This has significant implications for model providers – a cascade system is often made up of models which are not individually very powerful or risky – yet the overall system may exceed the FLOP threshold.
@@
FLOP varies dramatically for different modalities. In Figure 6(b), we plot the FLOP requirements over time of models grouped according to modality and downstream use case (model FLOP data from Epoch AI (2024)). It is easy to observe that the compute requirements have not increased at the same rate across modalities. For example, code models typically require less compute (Lin et al., 2024b), as do biological models (Maug et al., 2024). Multilingual models (Üstün et al., 2024; Aryabumi et al., 2024) tend to require more compute for each additional language covered. This is often referred to as the curse of multilinguality (Üstün et al., 2024; Arivazhagan et al., 2019; Conneau et al., 2019; Pfeiffer et al., 2022), where capacity is split between more languages such that performance on any given language suffers relative to a monolingual (single language) model of the same size. These differing compute needs mean that a single threshold may penalize some types of models and reward others. For example, thresholds may penalize multilingual models that attempt to serve many languages and improve access to technology (Üstün et al., 2024; Aryabumi et al., 2024).
@@
One way to address differences in modalities is to maintain different compute thresholds for each modality. While at first glance this is an attractive solution, it also imposes more technical overhead on governments who must correctly set a hard-coded benchmark for each modality. For example, it is interesting to note that the US Executive Order already has at least one modality-specific caveat to the compute thresholds by carving out a separate compute threshold for biological models. It is set lower for models trained for biological sequence data at $10^{23}$. However, since the threshold was set, models like xTrimoPGLM (Chen et al., 2024) already exceed the biological threshold set at $231e$ 23 operations by a factor of 6x (Maug et al., 2024). Many models (Lin et al., 2023; Elnaggar et al., 2020; Dalla-Torre et al., 2023) are currently within a factor of 10x the Executive Order’s reporting threshold (Maug et al., 2024). These models do not appear to present a decidedly different risk profile from previous generations, so if the goal of the thresholds is to be an inflection point for amplified risk it is unclear if it has been set successfully.
@@
Specifying separate thresholds for different modalities also risks inviting gamification. For example, to avoid a lower threshold for scrutiny for biological models one loophole is to preserve biology specific training data at less than 50%. According to current guidance the model would no-longer qualify as a “biological” model and would only be subject to the higher general purpose compute thresholds. Galactica-120B (Taylor et al., 2022) and Llama-molinst-protein-7b (Fang et al., 2024b) are both examples of models with capabilities for biological sequence modeling without primarily being trained on biological sequence data. Despite both presenting biological capabilities, neither is likely to be considered “biological” under the current Executive Order requirements (Maug et al., 2024). This highlights the fundamental tension of relying on compute alone – since it is not anchored to the risk metric that is of primary concern, it may be possible to sidestep in many creative ways while still presenting high-risk capabilities.
@@ 4 We are not very good at predicting the relationship between compute and risk
(a) Train Compute
(b) Train Dataset Size
(c) Train Time
(d) Model Size
Figure 7: The distribution of model attributes from models designated as notable AI systems by Epoch AI (Epoch AI, 2024). All of these properties are heavily skewed, with a non-normal distribution. Note the histogram axis is set to log scale. This skew and the rapidly changing nature of these properties over time, makes it hard to apply a hard-coded threshold (fixed number of FLOP) with confidence. For rapidly changing distributions, dynamic automatically adjusting thresholds have historically been more successful as a policy tool.
(a) Chest Circumference
(b) Blood Pressure
(c) Resting Heart Rate
(d) Baby Weights
Figure 8: Many natural phenomena follow reasonably close to a normal distribution, such as chest circumference (Quetelet, 1817), diastolic blood pressure (Musameh et al., 2017), resting heart rate (Quer et al., 2020), distribution of baby weights (Shen et al., 2014). Historically, hard coding thresholds work well with normal distributions because the data is well-behaved and predictable. For example, hospitals trigger protocols around ICUs for newborns based on birthweight (Cutland et al., 2017; Seri & Evans, 2008). There are successful examples of governments setting hard thresholds when they designate speed limits (US Department of Transportation, 2020) or limits for blood alcohol to determine drinking under the influence (World Health Organization, ).
@@ H.T. Pledge
Compute thresholds could be much improved by moving to dynamic instread of static thresholds An unpredictable relationship between compute and performance means that there will likely be false negatives when a hard threshold is set. That is, as smaller models become more performant, models which should be audited because of the risk they present avoid doing so because they fall underneath the threshold. Furthermore, it is likely that policymakers will constantly have to revisit and redefine a sensible threshold, which imposes technical overhead and creates issues with credibility.
@@
Given the large variance in compute FLOP across modalities, AI regulators should also look to the rich body of work on reference class forecasting (Baerenbold, 2023), where forecasts are only made relative to similar basket of goods. For example, if you wanted to predict how long it takes to read a history textbook, it is less informative to take the average reading time for all books in the world and likely more precise to restrict to similar history books. This is already done when setting property prices (takes into account local neighborhoods) and assessing risk on financial assets. In turn, policymakers could consider grouping models by whether they are general purpose in intent or domain-specialized (biological model for example). This should again be combined with additional metrics as FLOP is insufficient and be implemented as a dynamic threshold to avoid the technical overhead of continual adjustments of several hard coded thresholds.
Compute should not be used alone as a proxy for risk In 1928, the Soviet Union embarked on a set of 5-year plans where the government set specific targets for industrial output, agricultural production, and other economic indicators (Erlich, 1967). The metrics for success where defined almost entirely by the quantity of goods built, rather than the quality. This under-specification led to decades of commendable success in growth of production, but extremely low quality output which was often immediately discarded (Duda, 2023). In the same vein, a clear takeaway is that compute cannot be used as the only indicator of risk.
@@ 5.2 Parting Thoughts
Policymakers face a formidable task ahead of them. What is humbling and, at times, overwhelming to ponder is that computer science as a discipline is incredibly young – it has been a mere 68 years since the Dartmouth workshop where the term Artificial Intelligence was coined. Much remains to be discovered, and new tools will pose formidable risks and benefits. Perhaps one of the key takeaways of this essay, is that we must have necessary humility about our ability to predict the future. Compute thresholds are currently presented as a very rigid governance tools because of the emphasis on a single static number to tier risk. These types of estimates are prone to failure precisely because of how rapidly the landscape is changing. Instead, we should focus on flexible tools for monitoring risk that are not tied to static numbers. Furthermore, FLOP as a measure can be greatly improved by standardizing reporting and closing possible loopholes. In the previous Section 5.1, we discussed some of these recommendations. As to what comes next, the only certain thing is that something will come next. Perhaps fitting to conclude with a quote from Alan Turing “We can only see a short distance ahead, but we can see plenty there that needs to be done.”
@@ 6 Acknowledgments
Thank you to many of my wonderful colleagues and peers who took time to provide valuable feedback on earlier versions of this essay. I do not have much time to write or think deeply in isolation about a topic these days. Unfortunately, my time is increasingly spent helping others create breakthroughs. However, I have greatly enjoyed the small parcels of time I have spent on this essay wrestling with these ideas. This essay felt important to write because it requires grappling with several topics that are timely: the changing relationship we have with compute, how we navigate the risks introduced by the technology we have helped build and how science should inform policy. I have decided to release it as an essay that reflects the evolution of my thought process. This was more enjoyable for me writing down my thoughts, and was perhaps necessary for getting this to the finish line given other demands on my time. We are in an interesting time; it is rare to see research progress that is adopted overnight. Computer science ideas do not just resonate in conference halls anymore, but profoundly impact the world around us. This merits accountability, evidence and care as we navigate this impact.
Thanks for valuable feedback from several colleagues across several drafts of this essay (in no particular order): Usman Anwar, Neil Thompson, Sanmi Kojeyo, Helen Toner, Lennard Heim, Irene Solaiman, Shayne Longpre, Leshem Choshen, Sasha Luccioni, Stephen Casper, Jaime Sevilla, Nitarshan Rajkumar, Patrick Lewis, Aaron Courville, Nick Frosst, Rishi Bommasani, Gary Marcus, Thomas Diettrich, Margaret Jennings, Marzieh Fadaee, Ahmet Ustun, Aidan Peppin, Arash Ahmadian, Yoshua Bengio, Ivan Zhang, Markus Anderljung, Alexander Popper. Perhaps unusually, I regularly try and stress test ideas by seeking to understand the strongest counterarguments. I typically learn more from those who hold different viewpoints, and for this essay I have tried to invite input from colleagues with a varied set of stances on compute thresholds. No need to identify these worthy critics, but a huge thank you to everyone who engaged fully with this piece by providing very meaningful and rich feedback that greatly improved it. Many thanks to Aidan Peppin for additional valuable proofreading. An additional thanks to Linus Chui for visual input on the normal distribution plots in Figure 8. Many thanks to Shivalika Singh for putting together the associated website to make this essay more accessible to those beyond the academic community.

### v2

@@ 3.1.3 How to handle Mixture of Experts (MoEs) and classic ensembling?
3.1.4 FLOP only accounts for a single model, but does not capture risk of the overall system.
The emphasis on compute thresholds as an indicator of risk also implies that risk is the property of a single model rather than the system in which it is deployed. In the real-world, impact and risk are rarely attributable to a single model but are a facet of the entire system a model sits in and the way it interacts with its environment (Zaharia et al., 2024; Sculley et al., 2015; Jatho et al., 2023; Raji et al., 2020). Many real-world production systems are made up of cascading models where the final output is produced as a results of inputs being processed by multiple algorithms in sequence (Paleyes et al., 2022; Forum, 2023; Sculley et al., 2015; Shankar et al., 2022). There has yet to be guidance on whether the FLOP threshold is specific to a single model or whether all models that constitute an end-to-end system contribute to the final tally. This has significant implications for model providers – a cascade system is often made up of models which are not individually very powerful or risky – yet the overall system may exceed the FLOP threshold.
@@
3.1.5 FLOP varies dramatically across different modalities.
@@
In Figure 5(b), we plot the FLOP requirements over time of models grouped according to modality and downstream use case (model FLOP data from Epoch AI (2024)). It is easy to observe that the compute requirements have not increased at the same rate across modalities. For example, code models typically require less compute (Lin et al., 2024b), as do biological models (Maug et al., 2024). Multilingual models (Üstün et al., 2024; Aryabumi et al., 2024) tend to require more compute for each additional language covered. This is often referred to as the curse of multilinguality (Üstün et al., 2024; Arivazhagan et al., 2019; Conneau et al., 2019; Pfeiffer et al., 2022), where capacity is split between more languages such that performance on any given language suffers relative to a monolingual (single language) model of the same size. These differing compute needs mean that a single threshold may penalize some types of models and reward others. For example, thresholds may penalize multilingual models that attempt to serve many languages and improve access to technology (Üstün et al., 2024; Aryabumi et al., 2024).
One way to address differences in modalities is to maintain different compute thresholds for each modality. While at first glance this is an attractive solution, it also imposes more technical overhead on governments who must correctly set a hard-coded benchmark for each modality.
Figure 6: Bytes Magazine Cover, Volume 10, 1985. A key difficulty setting compute thresholds is that different domains and downstream tasks (language, vision, biology) demand very different levels of training compute, and so one compute threshold is not suitable to rule them all. This imposes more technical overhead on governments who must correctly set a hard-coded benchmark for each area. Only one domain specific compute threshold has been set to-date, by the EO for biological models. However, it has already been surpassed by several models that do not clearly present more risk than previous generations so may have been set too low.
@@
For example, it is interesting to note that the US Executive Order already has at least one modality-specific caveat to the compute thresholds by carving out a separate compute threshold for biological models. It is set lower for models trained for biological sequence data at $10^{23}$. However, since the threshold was set, models like xTrimoPGLM (Chen et al., 2024) already exceed the biological threshold set at $231e$ 23 operations by a factor of 6x (Maug et al., 2024). Many models (Lin et al., 2023; Elnaggar et al., 2020; Dalla-Torre et al., 2023) are currently within a factor of 10x the Executive Order’s reporting threshold (Maug et al., 2024). These models do not appear to present a decidedly different risk profile from previous generations, so if the goal of the thresholds is to be an inflection point for amplified risk it is unclear if it has been set successfully.
@@
Specifying separate thresholds for different modalities risks inviting gamification. For example, to avoid a lower threshold for scrutiny for biological models one loophole is to preserve biology specific training data at less than 50%. According to current guidance the model would no-longer qualify as a “biological” model and would only be subject to the higher general purpose compute thresholds. Galactica-120B (Taylor et al., 2022) and Llama-molinst-protein-7b (Fang et al., 2024b) are both examples of models with capabilities for biological sequence modeling without primarily being trained on biological sequence data. Despite both presenting biological capabilities, neither is likely to be considered “biological” under the current Executive Order requirements (Maug et al., 2024). This highlights the fundamental tension of relying on compute alone – since it is not anchored to the risk metric that is of primary concern, it may be possible to sidestep in many creative ways while still presenting high-risk capabilities.
@@ 5.1 Moving Away from Hard Coded Compute Thresholds
5.2 The case for dynamic thresolds.
H.T. Pledge
Compute thresholds could be much improved by moving to dynamic instread of static thresholds. An unpredictable relationship between compute and performance means that there will likely be false negatives when a hard threshold is set. That is, as smaller models become more performant, models which should be audited because of the risk they present avoid doing so because they fall underneath the threshold. Furthermore, it is likely that policymakers will constantly have to revisit and redefine a sensible threshold, which imposes technical overhead and creates issues with credibility.
@@ H.T. Pledge
(a) Compute
(b) Dataset Size
(c) Train Time
(d) Model Size
Figure 7: The distribution of model attributes from models designated as notable AI systems by Epoch AI (Epoch AI, 2024). All of these properties are heavily skewed, with a non-normal distribution. Note the histogram axis is set to log scale. This skew and the rapidly changing nature of these properties over time, makes it hard to apply a hard-coded threshold with confidence. For rapidly changing distributions, dynamic automatically adjusting thresholds have historically been more successful as a policy tool.
(a) Chest
(b) Blood Pressure
(c) Heart Rate
(d) Baby Weights
Figure 8: Many natural phenomena follow reasonably close to a normal distribution, such as chest circumference (Quetelet, 1817), diastolic blood pressure (Musameh et al., 2017), resting heart rate (Quer et al., 2020), distribution of baby weights (Shen et al., 2014). Historically, hard coded static thresholds work well with normal distributions because the data is well-behaved and predictable.
@@
Percentile based measures can also take into account differences in modality Given the large variance in compute FLOP across modalities, AI regulators should also look to the rich body of work on reference class forecasting (Baerenbold, 2023), where forecasts are only made relative to similar basket of goods. For example, if you wanted to predict how long it takes to read a history textbook, it is less informative to take the average reading time for all books in the world and likely more precise to restrict to similar history books. This is already done when setting property prices (takes into account local neighborhoods) and assessing risk on financial assets. In turn, policymakers could consider grouping models by whether they are general purpose in intent or domain-specialized (biological model for example). This should again be combined with additional metrics as FLOP is insufficient and be implemented as a dynamic threshold to avoid the technical overhead of continual adjustments of several hard coded thresholds.
5.3 Compute should not be used alone as a proxy for risk.
In 1928, the Soviet Union embarked on a set of 5-year plans where the government set specific targets for industrial output, agricultural production, and other economic indicators (Erlich, 1967). The metrics for success where defined almost entirely by the quantity of goods built, rather than the quality. This under-specification led to decades of commendable success in growth of production, but extremely low quality output which was often immediately discarded (Duda, 2023). In the same vein, a clear takeaway is that compute cannot be used as the only indicator of risk.
@@ 5.4 Parting Thoughts
Policymakers face a formidable task ahead of them. What is humbling and, at times, overwhelming to ponder is that computer science as a discipline is incredibly young – it has been a mere 68 years since the Dartmouth workshop where the term Artificial Intelligence was coined. Much remains to be discovered, and new tools will pose formidable risks and benefits. Perhaps one of the key takeaways of this essay, is that we must have necessary humility about our ability to predict the future. Compute thresholds are currently presented as a very rigid governance tool because of the emphasis on a single static number to tier risk. These types of estimates are prone to failure precisely because of how rapidly the landscape is changing. Instead, we should focus on flexible tools for monitoring risk that are not tied to static numbers. Furthermore, FLOP as a measure can be greatly improved by standardizing reporting and closing possible loopholes. In the previous Section 5.1, we discussed some of these recommendations. As to what comes next, the only certain thing is that something will come next. Perhaps fitting to conclude with a quote from Alan Turing “We can only see a short distance ahead, but we can see plenty there that needs to be done.”
@@ 6 Acknowledgments
Thank you to many of my wonderful colleagues and peers who took time to provide valuable feedback on earlier versions of this essay. I do not have much time to write or think deeply in isolation about a topic these days. My time is increasingly spent helping others create breakthroughs. However, I have greatly enjoyed the small parcels of time I have spent on this essay wrestling with these ideas. This essay felt important to write because it requires grappling with several topics that are timely: the changing relationship we have with compute, how we navigate the risks introduced by the technology we have helped build and how science should inform policy. We are in an interesting time; it is rare to see research progress that is adopted overnight. Computer science ideas do not just resonate in conference halls anymore, but profoundly impact the world around us. This merits accountability, evidence and care as we navigate this impact.
Thanks for valuable feedback from several colleagues across several drafts of this essay (in no particular order): Usman Anwar, Neil Thompson, Sanmi Koyejo, Helen Toner, Lennard Heim, Irene Solaiman, Shayne Longpre, Leshem Choshen, Sasha Luccioni, Stephen Casper, Jaime Sevilla, Nitarshan Rajkumar, Patrick Lewis, Aaron Courville, Nick Frosst, Rishi Bommasani, Gary Marcus, Thomas Diettrich, Margaret Jennings, Marzieh Fadaee, Ahmet Ustun, Aidan Peppin, Arash Ahmadian, Yoshua Bengio, Ivan Zhang, Markus Anderljung, Alexander Popper. Perhaps unusually, I regularly try and stress test ideas by seeking to understand the strongest counterarguments. I typically learn more from those who hold different viewpoints, and for this essay I have tried to invite input from colleagues with a varied set of stances on compute thresholds. No need to identify these worthy critics, but a huge thank you to everyone who engaged fully with this piece by providing very meaningful and rich feedback that greatly improved it. Many thanks to Aidan Peppin for additional valuable proofreading. An additional thanks to Linus Chui for visual input on the normal distribution plots in Figure 8. Many thanks to Shivalika Singh for putting together the associated website to make this essay more accessible to those beyond the academic community.